OpenStudy (anonymous):

find the zeros of each function state the multiplicity of multiple zeros y=x^2(x+4)^3(x-1)

OpenStudy (anonymous):

That's easy. All you have to do is set each factor to equal zero and go on from there.

OpenStudy (anonymous):

can you show step by step please i cant remember how to do this

OpenStudy (anonymous):

(x+4)=0... (x-1)=0. that's the first step. The rest should be clearer now.